[PIPE2D-1473] bad flux calibration in GE visits Created: 14/Jun/24  Updated: 17/Feb/25  Resolved: 17/Feb/25

Status: Done
Project: DRP 2-D Pipeline
Component/s: None
Affects Version/s: None
Fix Version/s: None

Type: Task Priority: Normal
Reporter: Masayuki Tanaka Assignee: Unassigned
Resolution: Done Votes: 0
Labels: EngRun, flux-calibration
Remaining Estimate: Not Specified
Time Spent: Not Specified
Original Estimate: Not Specified

Attachments: PNG File fluxcal_fix_after.png     PNG File fluxcal_fix_before.png     PNG File fluxcal_many_visits.png     PNG File tmp - 2024-06-10T220137.058.png    
Issue Links:
Relates
relates to PIPE2D-1476 Fix n2 detectorMap shifted by 1 fiber Done
Sprint: postRun17, preRun18August1, EngRun18August, PostRun18, preRun19Oct2024, postRun19Oct, preRun20Jan

 Description   

rhl pointed out the flux calibration is catastrophically bad in GE visits. I am filing this ticket as per Mortani-san's request.

 



 Comments   
Comment by Masayuki Tanaka [ 14/Jun/24 ]

I confirm the flux calibration is indeed bad. I picked visit=111775 from the GE visits and the flux calibration vector looks awful. This seems to come from bad pfsMerged in spectrograph=2. Robert pointed out that n2 has inconsistent fluxes between real data and targetType. I confirm that it is only n2 and the other arms+spectrographs are OK. This seems to be a detectorMap problem.
-----------------------------------

#fiberId targetType b2    r2      n2
658 FLUXSTD 2572.8 5661.3 1606.0
659 FLUXSTD 1805.8 4226.6 405.5
660 SKY 53.7 116.2 494.0
661 SKY 52.1 116.8 1791.0
662 FLUXSTD 1539.3 3674.8 459.9
663 SKY 45.2 112.0 451.9
664 SKY 53.6 124.6 429.4
665 SKY 44.2 114.1 1393.1
666 FLUXSTD 1227.6 2884.7 529.0
667 SKY 39.7 105.4 499.6

-----------------------------------

The latter 3 columns are the median flux in each arm and is from pfsArm.

Comment by Kiyoto Yabe [ 16/Jul/24 ]

Although a complete check is not yet, here is an example of flux calibrated spectrum (a FLUXSTD) before and after the fix. I looks the calibration is significantly improved.

 

Left: before the fix, Right: after the fix

Multiple visits for the same object.

Comment by Kiyoto Yabe [ 15/Feb/25 ]

Masayuki Tanaka Are we going to keep investigating the flux loss issue (of about 20-30%) in this ticket? or would PIPE2D-1580 be a better place?

Comment by Masayuki Tanaka [ 17/Feb/25 ]

This ticket is about the bad fluxCal due to off-by-one problem in the detectorMap. So, we can safely close this ticket. For the 20-40% flux loss issue, I will file a separate ticket. There is also the nearly-no-flux problem but we can track that work on PIPE2D-1580.

Generated at Sun Apr 06 09:50:18 JST 2025 using Jira 8.3.4#803005-sha1:1f96e09b3c60279a408a2ae47be3c745f571388b.